OpenStudy (anonymous):

Use rhombus TQRS below for questions 1–4.

OpenStudy (anonymous):

1. What is the measure of angle 1 ? A. 47 B. 37 C. 74 D. 53

OpenStudy (anonymous):

2. What is the measure of angle 2 ? A. 47 B. 74 C. 37 D. 53

ganeshie8 (ganeshie8):

one q at a time please

OpenStudy (anonymous):

1. What is the measure of angle 1 ? A. 47 B. 37 C. 74 D. 53

ganeshie8 (ganeshie8):

lol ok, use this for q 1 :- in rhombus, diagonals bisect the angles

OpenStudy (anonymous):

I'm confused

ganeshie8 (ganeshie8):

what does that mean ? the diagonal SQ divides the angle S into two equal halves

ganeshie8 (ganeshie8):

look at pix, so angle 1, and 53 are same

ganeshie8 (ganeshie8):

angle 1 = 53

OpenStudy (anonymous):

ok take a look at q 2 please

ganeshie8 (ganeshie8):

ok, use this for q 2 :- in rhombus, consecutive angles add up to 180

ganeshie8 (ganeshie8):

S and T are consecutive angles here

OpenStudy (anonymous):

so do you do 180-53 ?

ganeshie8 (ganeshie8):

kinda, but 53 is not full S

ganeshie8 (ganeshie8):

53 is only half of S, so we get this :- T+S = 180 T + 53*2 = 180 T + 106 = 180 T = 74 => angle 2 = T/2 = 74/2 = ?

OpenStudy (anonymous):

37 ?

OpenStudy (anonymous):

Correct! :D